OpenStudy (anonymous):

If the costs of goods sold was $42,000, the income from sales was $86,000, and the expenses were $12,000, find (50) gross profit and (51) net profit.

OpenStudy (campbell_st):

Gross profit = sales - costs = 86000 - 42000 = 44000 Net profit = Gross Profit - Expenses = 44000 - 12000

OpenStudy (anonymous):

Thank you! :)